OpenStudy (anonymous):

solve for w pleaseeee !!! 8=7/w-8

OpenStudy (homeworkhelper099):

Hi phm96, please explain to me which math you are taking.

OpenStudy (anonymous):

this is pre algbra

OpenStudy (misty1212):

HI!!

OpenStudy (amorfide):

multiply both sides by the bottom of the fraction to create your magic number 1 8(w-8)=7(w-8)/(w-8) the w-8 will cancel to 1 to get 8(w-8)=7(1)

OpenStudy (misty1212):

is this \[8=\frac{7}{w-8}\]?

OpenStudy (amorfide):

you can finish the rest i am sure

OpenStudy (anonymous):

multiply both sides by (w-8)

OpenStudy (anonymous):

then u get \((w-8)\times 8=7\)

OpenStudy (anonymous):

simplify \(8w-64=7\)

OpenStudy (anonymous):

add 64 to both sides \(8w=71\)

OpenStudy (anonymous):

divide both sides by 8 \[w=8 \frac{ 7 }{ 8 }\]

OpenStudy (anonymous):

is that what you are looking for?

OpenStudy (anonymous):

tell me if u don't understand